OpenStudy (anonymous):

Rewrite with only sin x and cos x sin 3x + cos 3x

OpenStudy (amistre64):

sin(3x) = sin(2x + x) = sin(2x)cos(x) + sin(x)cos(2x) = [sin(x)cos(x) + sin(x)cos(x)]cos(x) + sin(x)[cos(x)cos(x) - sin(x)sin(x)] ------------------------------------------ cos(3x) = cos(2x+x) = cos(2x)cos(x)-sin(2x)sin(x) = [cos(x)cos(x)-sin(x)sin(x)]cos(x) - [sin(x)cos(x)+sin(x)cos(x)]sin(x)

OpenStudy (amistre64):

sin(3x) + cos(3x) = 2 sin(x)cos^2(x) + cos^2(x)sin(x) - sin^3(x) + cos^3(x) -sin^2(x) cos(x) - sin^2(x) cos(x) - sin^2(x) cos(x) simplify as needed